OpenStudy (anonymous):

10 ^ 3x-8 =2^5-x This section is giving me major problems.

OpenStudy (anonymous):

\[\huge 10^{3x-8}=2^{5-x}\] here's the trick. Always make all your terms into prime numbers.

OpenStudy (anonymous):

10 is not a prime number. So we need to split it into two numbers.

OpenStudy (anonymous):

So we take a factor of 10 that is a prime number. What number can you think of that is a factor of 10 but also a prime number?

OpenStudy (anonymous):

2 and 5

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Now we take a property of indices where \[\large (ab)^2=a^2b^2\]

OpenStudy (anonymous):

i dont understand how that helps

OpenStudy (anonymous):

So what would the LHS look like after you've just learnt that you need to split the numbers and then this property?

OpenStudy (anonymous):

Okay...

OpenStudy (anonymous):

Let's split it first shall well?

OpenStudy (anonymous):

\[\large (5\times 2)^{3x-8}=2^{5-x}\]

OpenStudy (anonymous):

get rid of the brackets using that property I told you.

OpenStudy (anonymous):

Let me just give you the property again without the power of 2. \[\large (a\times b)^n=a^nb^n\]

OpenStudy (anonymous):

5^3x-8 2^5-x

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

2^3x-8

OpenStudy (anonymous):

Well Done mate. Fantastic.

OpenStudy (anonymous):

Now this next part is like collecting like terms.

OpenStudy (anonymous):

Have you heard about collecting like terms before?

OpenStudy (anonymous):

i dont know how to do that

OpenStudy (anonymous):

So you haven't heard about colelcting like terms when you did algebra. Collecting like terms: x+x+y+2y=2x+3y?

OpenStudy (anonymous):

i dont have a clue with this mess

OpenStudy (anonymous):

Okay... \[\huge 5^{3x-8}2^{3x-8}=2^{5-x}\] \[\huge 5^{3x-8}=\frac{2^{5-x}}{2^{3x-8}}\]

OpenStudy (anonymous):

Simplify the RHS (Right Hand Side) for me please.

OpenStudy (anonymous):

is it 5-x/3x-8?

OpenStudy (anonymous):

\[\huge \frac{a^{n-1}}{a^{n+3}}=a^{(n-1)-(n+3)}\]

OpenStudy (anonymous):

oooo

OpenStudy (anonymous):

2^(5-x)-(3x-8)

OpenStudy (anonymous):

Correct. Now simplify the indice.

OpenStudy (anonymous):

2^-3-4x

OpenStudy (anonymous):

\[5-x-(3x-8)\neq 5-x-3x-8\] \[5-x-(3x-8)=5-x-3x--8\] \[5-x-3x--8=5-x-3x+8\]

OpenStudy (anonymous):

is that the final answer?

OpenStudy (anonymous):

No, that was simlifying that indice of yours. You can't seem to understand that two minus signs become a plus sign....

OpenStudy (anonymous):

simplifying*

OpenStudy (anonymous):

i understand that

OpenStudy (anonymous):

So what's the simplified form of \[\huge 2^{(5-x)-(3x-8)}\]

OpenStudy (anonymous):

is it 13 -4x?

OpenStudy (anonymous):

Great!! Now you're starting to get it.

OpenStudy (anonymous):

lol i am dumb blonde

OpenStudy (anonymous):

Okay. Next step. We just simplified the RHS. \[\huge 5^{3x-8}=2^{13-4x}\]

OpenStudy (anonymous):

but i am just a sophmore

OpenStudy (anonymous):

that means nothing to me

OpenStudy (anonymous):

That still doesn't make the cut of doing good math. I was in a grade less than that, and I was still excelling at this stuff.

OpenStudy (anonymous):

Okay we convert 5^3x-8 into log.

OpenStudy (anonymous):

yikes

OpenStudy (anonymous):

is it logbase5 (3x-8)

OpenStudy (anonymous):

sorry I meant convert 2^13-4x into log. My apologies.

OpenStudy (anonymous):

is it log 2 13-4x

OpenStudy (anonymous):

Here's what you should be doing to convert it. \[\huge a^n=b^x\] \[\huge log_{b}a^{n}=x\]

OpenStudy (anonymous):

i am new to this and i dont have a clue

OpenStudy (anonymous):

There's nothing new about maths. It doesn't matter if you're new to this website or not. Being new to the website doesn't necessarily mean you're new to math...

OpenStudy (anonymous):

I am new to logs this week, and I have the worst math teacher in the school, I do not understand logs yet, I am sorry. I am normally a very good student.

OpenStudy (anonymous):

You're meant to know the laws of logarithms before doing this. Is the correct answer in decimal places?

OpenStudy (anonymous):

Because this questions is seriously a messy one. I need to round off to get the answer.

OpenStudy (anonymous):

this question*

OpenStudy (anonymous):

you are very smart, and this is my first week with logarithms

OpenStudy (anonymous):

i will get them, in a few days it will be easy for me, but right now i am completely lost

OpenStudy (anonymous):

I wanted to ask whether you have the correct answer on an answer sheet? If you do is the correct answer in decimal places?

OpenStudy (anonymous):

I do not have any answer to the problem.

OpenStudy (anonymous):

Okay. we'll continue. What did you get when you convert the terms into log. Look at that property I gave you.

OpenStudy (anonymous):

I dont know how to do it

OpenStudy (anonymous):

Here's a straight example. Apply this example to our current method okay? \[\huge x^{n+1}=y^{n-2}\] \[\huge log_{x}y^{n-2}=n+1\]

OpenStudy (anonymous):

current question**

OpenStudy (anonymous):

I have 2^-4x+13

OpenStudy (anonymous):

You make 2 the log base.

OpenStudy (anonymous):

the log base is the n in this: \[\huge log_{n}<---log base\]

OpenStudy (anonymous):

log base 2

OpenStudy (anonymous):

Yes. Now the number beside log base 2 is the WHOLE TERM on the LHS.

OpenStudy (anonymous):

THe whole term is \[\huge 5^{3x-8}\]

OpenStudy (anonymous):

log base 2 5^3x-8 = -4x +13

OpenStudy (anonymous):

Good work.

OpenStudy (anonymous):

Now we need to apply another property of logs... \[\huge log_{n}x^a=alog_{n}x\]

OpenStudy (anonymous):

i thought we were done

OpenStudy (anonymous):

The a is the x^a can be moved to the side of the log. And no, we aren't done. You're meant to solve for x, aren't you?

OpenStudy (anonymous):

the a in*

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

3x-8 log base 2 5

OpenStudy (anonymous):

what? this is what you had: log base 2 5^3x-8 = -4x +13?

OpenStudy (anonymous):

And yes you're correct with the LHS.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

does the right stay the same

OpenStudy (anonymous):

\[\huge (3x-8)log_{2}5=13-4x\]

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Move the (3x-8) to the RHS.

OpenStudy (anonymous):

\[\huge log_{2}5=\frac{13-4x}{3x-8}\]

OpenStudy (anonymous):

yes i have that

OpenStudy (anonymous):

Then, we get rid of the logs. We will get this. \[\huge 2^{\frac{13-4x}{3x-8}}=5\]

OpenStudy (anonymous):

to find what the indice of 2 is equal to, we must use logs again. So... \[\frac{13-4x}{3x-8}=\frac{log5}{log2}\] You should be able to find what the RHS is equal to when you plug that into your calculator.

OpenStudy (anonymous):

Thank you very much, I have to go to bed, i am up hours past my bed time

OpenStudy (anonymous):

I will be back and hope I see you again

OpenStudy (anonymous):

Well you can probably find x by yourself after that.

OpenStudy (anonymous):

And no worries.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Good night.

OpenStudy (anonymous):

you too